A rectangle of area 350 square feet is 14 times as wide as it is long. Find its length and width
Answer by Cromlix(4381)

Hi there,
Make the length = 'x'
Width = 14x
Area of rectangle = length x width
350 = x(14x)
14x^2 = 350
x^2 = 25
x = 5ft
Width = 14(5) = 70ft
Length = 5ft
Hope this helps:-)